OpenStudy (anonymous):

tennisball mashine shoots a ball straight up in the air, it lands after 7 seconds What's the hight of the ball given v=f(t)=30-10t i figured that the equation i got is actually the derivative of -5t^2+30t ball shots upp with a velocity of 30m/s, is a physics mix i have no pheta angle neede to solve it in a simpler way...vertex apex im asked: what does 30 stand for? -10t?acceleration? and what's max height? i have a graf with and dowsslopping linear line crossing the y axis at 30m/s and the x axis at 3 seconds

OpenStudy (anonymous):

ok tell me V=30-10t did they give you this with question

OpenStudy (anonymous):

oh ok now i can solve it

OpenStudy (anonymous):

is the answer 35 feet

OpenStudy (anonymous):

because you just sub 7 into original equation and you will get maximum height

OpenStudy (anonymous):

\[=-5(7)^{2}+30 \times 7\]

OpenStudy (irishboy123):

yes, for \(v(t) =30 - 10t\) we have 30 as initial velocity, -10 for gravity (acceleration) so as you say \(x(t) =30t - 5t^2+x_0 = 5t(6-t)+x_0\) but note the integration constant if it lands after t = 7, then \(\large x(7) = x_0-35\) !!!!! weird! is the machine 35m above gound or does the ball land in a crater?!?! anyways, at the peak of its flight \(v(t) = 0 \implies 30 - 10t=0, t = 3\) solve from there